# Crossed product splitting of intermediate operator algebras via $2$-cocycles

### v1

@@ 1. Introduction
Operator algebra theory is not exception of this general philosophy. Studies of subalgebras (e.g., maximal amenable subalgebras [48]) and embeddings (e.g., injective enveloping C-algebras [22], constructions and classification of embeddings) are interested and deeply studied by a wide range of operator algebraists, and some of them become key players of major breakthroughs. Highlights include Jones’s subfactor theory [28], Popa’s deformation/rigidity theory [50], accessible characterizations on C-simplicity by Kalantar–Kennedy [29] and their striking consequences [8], the non-commutative Nevos–Zimmer theorem of Boutonnet–Houdayer [7], [23], and an abstract approach to the classification theorem of unital simple separable nuclear C-algebras [12].
@@
Let $\Gamma$ be a discrete group with Haagerup–Kraus’s approximation property ((((AP)))). Let $B\subset(A,\alpha)$ be a C-irreducible, centrally $\Gamma$-free inclusion of unital $\Gamma$-C-algebras. Then for any $C\in{\rm Int}(B\subset A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma)$, there is a subgroup $\Lambda<\Gamma$ and an $(\mathcal{O}_{2}\otimes C,\Lambda)$-respecting map $\mathfrak{u}\colon\Lambda\rightarrow(\mathcal{O}_{2}\otimes A)^{\rm u}$ with
@@
There is an inclusion $B\subset(A,\alpha)$ as in Theorem A with the following property. There is $C\in{\rm Int}(B\subset A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma)$ such that for any unital C-algebra $D$ with $[1_{D}]_{0}\neq 0$, the tensor product $C\otimes D$ does not enjoy a natural twisted crossed product splitting.
@@
We also establish the von Neumann algebra version of Theorem A, where $\mathcal{O}_{2}$ is replaced by a type I factor (Corollary 7.7). Because the K0-obstruction is very small in factors, under assumptions on the Murray–von Neumann type, we do not need a tensor product component.
@@
    If $N$ is of infinite type, then for any $C\in{\rm Int}(N\subset M\mathop{{\bar{\rtimes}}_{\alpha}}\Gamma)$, there is a subgroup $\Lambda<\Gamma$ and an $\alpha$-$1$-cocycle $\mathfrak{u}\colon\Lambda\rightarrow M^{\rm u}$ with
@@ 2.1. Notations
  • For a C-algebra $A$, denote by $A^{\rm p}$ the set of all projections in $A$. Denote by $\mathcal{M}(A)$ the multiplier algebra of $A$. When $A$ is unital, we write $A^{\rm u}$ for the unitary group of $A$.
@@
  • We use the symbol ‘$<$’ for (not necessary proper) closed subgroups. So $H<G$ means that $H$ is a closed subgroup of a topological group $G$.
@@ 2.5. Twisted crossed product C-subalgebras in crossed products
given by sending $cs^{{\rm d}\mathfrak{u}}$ to $c\mathfrak{v}_{s}$ for $c\in C$ and $s\in\Lambda$. The canonical conditional expectation ${\mathbf{E}}\colon A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma\rightarrow A$ restricts to a faithful conditional expectation from the closure of the image of $\iota\rtimes\mathfrak{v}$ onto $C$. Moreover the restriction is an extension of the evaluation map ${\rm ev}_{e}$ on (the image of) $C\rtimes_{{\rm alg},{}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u}}\Lambda$. Thus $\iota\rtimes\mathfrak{v}$ extends to an embedding $C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u}}}\Lambda\rightarrow A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma$. Via this embedding, we identify $C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u}}}\Lambda$ with a C-subalgebra of $A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma$. We next see that this subalgebra (but the embedding) depends only on $({}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u})$ (not on the choice of $\mathfrak{u}$). This justifies the notation $C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u}}}\Lambda\subset A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma$. In fact we observe a slightly more general result.
@@ 3. Proof of Theorem A
We say a C-algebra inclusion $B\subset A$ is $\mathcal{O}_{2}$-absorbing if all members $C$ of ${\rm Int}(B\subset A)$ satisfies $C\cong\mathcal{O}_{2}\otimes C$.
@@
Let $s\in\Gamma$ be an element with ${\mathbf{E}}_{s}(C)\neq\{0\}$. We first show that ${\mathbf{E}}_{s}(C)\cap A^{\rm u}\neq\emptyset$. Take $c\in C$ with ${\mathbf{E}}_{s}(c)\neq 0$. By Lemma 3.1, one has ${\mathbf{E}}_{s}(c)s\in C$. Hence ${\mathbf{E}}_{s}(c){\mathbf{E}}_{s}(c)^{\ast}\in C\cap A$. By the assumptions on $B\subset A$, every $D\in{\rm Int}(B\subset A)$ is purely infinite simple with K$(D)=0$ 0. Hence one can find $x\in C\cap A$ satisfying
@@
This implies that $p:={\mathbf{E}}_{s}(c)^{\ast}x^{\ast}x{\mathbf{E}}_{s}(c)$ is a projection. Because $x{\mathbf{E}}_{s}(c)s\in C$, one has $p\in\alpha_{s}(C\cap A)$. By [15], one can find $v\in\alpha_{s}(C\cap A)$ with $v^{\ast}v=1$, $vv^{\ast}=p$. Then one has $x{\mathbf{E}}_{s}(c)vs=x{\mathbf{E}}_{s}(c)s\alpha_{s}^{-1}(v)\in C$. Hence $x{\mathbf{E}}_{s}(c)v\in{\mathbf{E}}_{s}(C)\cap A^{\rm u}$.
@@ 4. Twisted and/or non-unital case
$C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{v}\mathfrak{u}^{\ast}}(^{\mathfrak{u}}\alpha),{}^{\mathfrak{v}\mathfrak{u}^{\ast}}({}^{\mathfrak{u}}\mathfrak{w})}}\Lambda\subset A\mathop{\rtimes_{{\mathrm{r}},{{}^{\mathfrak{u}}\alpha},{}^{\mathfrak{u}}\mathfrak{w}}}\Gamma\quad{\rm~{}and~{}}\quad C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{v}}\alpha,{}^{\mathfrak{v}}\mathfrak{w}}}\Lambda\subset A\mathop{\rtimes_{{\mathrm{r}},\alpha,\mathfrak{w}}}\Gamma$
@@
$\theta(C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{v}}\alpha,{}^{\mathfrak{v}}\mathfrak{w}}}\Lambda)=C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{v}\mathfrak{u}^{\ast}}({{}^{\mathfrak{u}}\alpha}),{}^{\mathfrak{v}\mathfrak{u}^{\ast}}({}^{\mathfrak{u}}\mathfrak{w})}}\Lambda$
@@
Let $(\alpha,\mathfrak{w})\colon\Gamma\curvearrowright A$ be a cocycle action and let $B\subset(A,\alpha,\mathfrak{w})$. Then there is a map $\mathfrak{u}\colon\Gamma\rightarrow\mathcal{M}(B\otimes\mathbb{K})^{\rm u}$ satisfying the relation (${}^{\mathfrak{u}}(\mathfrak{w}\otimes 1_{\mathbb{K}})=1$.
@@
We observe that the proof of Lemma 3.1 works in the present situation, after replacing $s$, $ts$ appearing in the series expansions of $\tilde{c}_{i}$ therein by $s^{\mathfrak{w}}$, $t^{\mathfrak{w}}s^{\mathfrak{w}}$ (not $(ts)^{\mathfrak{w}}$) respectively. Then Lemma 3.3 also works in this setting. By using Lemma 4.2 instead of Lemma 2.2, the proof is complete. ∎
For any simple C-algebra $A$, the tensor product $\mathcal{O}_{2}\otimes A$ is purely infinite simple. In particular it contains nonzero projections, and all nonzero projections are mutually Murray–von Neumann equivalent. Moreover, when $A$ is non-unital and separable, for any nonzero $p\in(\mathcal{O}_{2}\otimes A)^{\rm p}$, one can find a sequence $(v_{n})_{n\in\mathbb{N}}$ in $\mathcal{O}_{2}\otimes A$ with
@@ 5. Cocycle actions which are very hard to untwist
satisfying the same axioms as the discrete group case. Here, for a separable C-algebra $A$, we equip ${\rm Aut}(A)$, $A$, $\mathcal{M}(A)^{\rm u}$ with the point-norm topology, norm topology, strict topology, respectively. Note that all these topologies are Polishable, hence they are standard as Borel spaces. Similarly, we require Borel measurability for the maps witnessing exterior equivalence.
It is fundamental that we only require Borel measurability of the maps. Indeed, for two main motivating examples, namely cocycle actions arising from group extensions and corners of ((((cocycle)))) actions, the resulting maps are typically discontinuous. However we note that, when $\alpha$ is a homomorphism, it is automatically continuous, because $\operatorname{\mathop{Aut}}(A)$ is a Polish group by [63], Theorem D.11. Similarly, for ordinary actions $\alpha\colon G\curvearrowright A$, any Borel $1$-cocycle $\mathfrak{u}$ of $\alpha$ is automatically strictly continuous. Indeed the map $G\ni s\mapsto\mathfrak{u}_{s}s\in\mathcal{M}(A\mathop{\rtimes_{{\mathrm{r}},\alpha}}G)^{\rm u}$ defines a Borel homomorphism by the cocycle relation, and hence it must be continuous in the strict topology again by [63], Theorem D.11. This proves the strict continuity of $\mathfrak{u}$. As a result, amenability of C-dynamical systems (see [2], [42]) is preserved by exterior equivalence.
@@ 5.1. Corner cocycle actions
For an lcsc group $G$, let $(\alpha,\mathfrak{w})\colon G\curvearrowright A$ be a cocycle action on a separable C-algebra $A$. Assume that $p\in A^{\rm p}$ satisfies the following condition: the projections $\alpha_{g}(p)$; $g\in G$ are mutually Murray–von Neumann equivalent. Then one can define the corner cocycle action $(\alpha^{p},\mathfrak{w}^{p})\colon G\curvearrowright pAp$ as follows. We first observe the following lemma.
@@
Let $G$ be a non-exact lcsc group. Then there is a cocycle action $(\alpha,\mathfrak{w})\colon G\curvearrowright\mathcal{O}_{2}$ with the following property. For any cocycle action $(\beta,\mathfrak{x})\colon G\curvearrowright A$ on a unital separable C-algebra, the $(\alpha\otimes\beta)$-$2$-cocycle $\mathfrak{w}\otimes\mathfrak{x}$ is not a coboundary.
@@
Let $(\beta,\mathfrak{x})\colon G\curvearrowright A$ be any cocycle action on a unital separable C-algebra $A$. Suppose that $(\alpha\otimes\beta,\mathfrak{w}\otimes\mathfrak{x})$ is exterior equivalent to a genuine action $\gamma$. Then $(\alpha\otimes\beta,\mathfrak{w}\otimes\mathfrak{x})\otimes 1_{\mathbb{K}}$ is exterior equivalent to $\gamma\otimes 1_{\mathbb{K}}$. By Lemma 5.3, $(\alpha,\mathfrak{w})\otimes 1_{\mathbb{K}}$ is exterior equivalent to $\sigma$. Also $(\beta,\mathfrak{x})\otimes 1_{\mathbb{K}}$ is exterior equivalent to a genuine action by the stabilization trick [55], [44]. Since $\mathbb{K}\cong\mathbb{K}\otimes\mathbb{K}$, these observations show that $\gamma\otimes 1_{\mathbb{K}}$ is exterior equivalent to an amenable action. This implies the amenability of $\gamma$. Since $G$ is non-exact, this contradicts to Corollary 3.6 of [42]. ∎
@@
    When $G$ is exact, such a cocycle action does not exist. Indeed by combining the stabilization trick [55], [44] and the Gabe–Szabo classification theorem [18], one concludes the following statement: Let $\beta\colon G\curvearrowright\mathcal{O}_{2}$ be an amenable, isometrically shift-absorbing action [18]. Then, for any cocycle action $(\alpha,\mathfrak{w})\colon G\curvearrowright A$ on a simple separable nuclear C-algebra (in particular when $A=\mathcal{O}_{2}$), the $(\alpha\otimes\beta)$-$2$-cocycle $\mathfrak{w}\otimes 1$ is a coboundary.
@@
    Sutherland [55] proved that any $2$-cocycle of a cocycle action on a properly infinite von Neumann algebra is a coboundary. Thus Theorem 5.4 is a very unique phenomenon in C-algebra theory.
@@
6. K0-obsrtuction: New tensor component $\mathcal{O}_{2}$ is necessary and minimal

### v2

@@ 1. Introduction
Operator algebra theory is no exception to this general philosophy. Studies of subalgebras (e.g., maximal amenable subalgebras [48]) and embeddings (e.g., injective enveloping C-algebras [22], constructions and classification of embeddings) are of great interest and deeply studied by a wide range of operator algebraists, and some of them become key players of major breakthroughs. Highlights include Jones’s subfactor theory [28], Popa’s deformation/rigidity theory [50], accessible characterizations on C-simplicity by Kalantar–Kennedy [29] and their striking consequences [8], the non-commutative Nevos–Zimmer theorem of Boutonnet–Houdayer [7], [23], and an abstract approach to the classification theorem of unital simple separable nuclear C-algebras [12].
@@
Let $\Gamma$ be a discrete group with Haagerup–Kraus’s approximation property ((AP)). Let $B\subset(A,\alpha)$ be a C-irreducible, centrally $\Gamma$-free inclusion of unital $\Gamma$-C-algebras. Then for any $C\in{\rm Int}(B\subset A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma)$, there is a subgroup $\Lambda<\Gamma$ and an $(\mathcal{O}_{2}\otimes C,\Lambda)$-respecting map $\mathfrak{u}\colon\Lambda\rightarrow(\mathcal{O}_{2}\otimes A)^{\rm u}$ with
@@
There is an inclusion $B\subset(A,\alpha)$ as in Theorem A with the following property. There is $C\in{\rm Int}(B\subset A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma)$ such that for any unital C-algebra $D$ with $1_{D}]_{0}\neq 0$, the tensor product $C\otimes D$ does not enjoy a natural twisted crossed product splitting.
@@
We also establish the von Neumann algebra version of Theorem A, where $\mathcal{O}_{2}$ is replaced by a type I factor (Corollary 7.7). Because the K0-obstruction is very small in factors, under assumptions on the Murray–von Neumann type, we do not need to add a tensor product component.
@@
    If $N$ is of infinite type, then for any $C\in{\rm Int}(N\subset M\mathop{{\bar{\rtimes}}_{\alpha}}\Gamma)$, there is a subgroup $\Lambda<\Gamma$ and an $\alpha$-11-cocycle $\mathfrak{u}\colon\Lambda\rightarrow M^{\rm u}$ with
@@ 2.1. Notations
  • For a C-algebra AA, denote by $A^{\rm p}$ the set of all projections in $A$. Denote by $\mathcal{M}(A)$ the multiplier algebra of $A$. When $A$ is unital, we write $A^{\rm u}$ for the unitary group of $A$.
@@
  • We use the symbol ‘<<’ for (not necessary proper) closed subgroups. So $H<G$ means that $H$ is a closed subgroup of a topological group $G$.
@@ 2.5. Twisted crossed product C-subalgebras in crossed products
given by sending $cs^{{\rm d}\mathfrak{u}}$ to $c\mathfrak{v}_{s}$ for $c\in C$ and $s\in\Lambda$. The canonical conditional expectation ${\mathbf{E}}\colon A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma\rightarrow A$ restricts to a faithful conditional expectation from the closure of the image of $\iota\rtimes\mathfrak{v}$ onto $C$. Moreover the restriction is an extension of the evaluation map ${\rm ev}_{e}$ on (the image of) $C\rtimes_{{\rm alg},{}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u}}\Lambda$. Thus $\iota\rtimes\mathfrak{v}$ extends to an embedding $C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u}}}\Lambda\rightarrow A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma$. Via this embedding, we identify $C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u}}}\Lambda$ with a C-subalgebra of $A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma$. We next see that this subalgebra depends only on $({}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u})$ (not on the choice of $\mathfrak{u}$). This justifies the notation $C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{u}}\alpha,{\rm d}\mathfrak{u}}}\Lambda\subset A\mathop{\rtimes_{{\mathrm{r}},\alpha}}\Gamma$. In fact we observe a slightly more general result.
@@ 3. Proof of Theorem A
We say a C-algebra inclusion $B\subset A$ is $\mathcal{O}_{2}$-absorbing if all members $C$ of ${\rm Int}(B\subset A)$ satisfy $C\cong\mathcal{O}_{2}\otimes C$.
@@
[…]
@@
This implies that $p:={\mathbf{E}}_{s}(c)^{\ast}x^{\ast}x{\mathbf{E}}_{s}(c)$ is a projection. Because $x{\mathbf{E}}_{s}(c)s\in C$, one has $p\in\alpha_{s}(C\cap A)$. By [15], one can find $v\in\alpha_{s}(C\cap A)$ with $^{\ast}v=1, vv$$v^{\ast}=p$. Then one has $x{\mathbf{E}}_{s}(c)vs=x{\mathbf{E}}_{s}(c)s\alpha_{s}^{-1}(v)\in C$. Hence $x{\mathbf{E}}_{s}(c)v\in{\mathbf{E}}_{s}(C)\cap A^{\rm u}$.
@@
This partial splitting still holds true without C-irreducibility, but in this case, the subset $\Lambda$ needs not form a subgroup.
@@ 4. Twisted and/or non-unital case
$C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{v}\mathfrak{u}^{\ast}}\hskip-2.84526pt(^{\mathfrak{u}}\alpha),{}^{\mathfrak{v}\mathfrak{u}^{\ast}}\hskip-2.84526pt({}^{\mathfrak{u}}\mathfrak{w})}}\Lambda\subset A\mathop{\rtimes_{{\mathrm{r}},{{}^{\mathfrak{u}}\alpha},{}^{\mathfrak{u}}\mathfrak{w}}}\Gamma\quad{\rm~and~}\quad C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{v}}\alpha,{}^{\mathfrak{v}}\mathfrak{w}}}\Lambda\subset A\mathop{\rtimes_{{\mathrm{r}},\alpha,\mathfrak{w}}}\Gamma$
@@
$\theta(C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{v}}\alpha,{}^{\mathfrak{v}}\mathfrak{w}}}\Lambda)=C\mathop{\rtimes_{{\mathrm{r}},{}^{\mathfrak{v}\mathfrak{u}^{\ast}}\hskip-2.84526pt({{}^{\mathfrak{u}}\alpha}),{}^{\mathfrak{v}\mathfrak{u}^{\ast}}\hskip-2.84526pt({}^{\mathfrak{u}}\mathfrak{w})}}\Lambda$
@@
Let $(\alpha,\mathfrak{w})\colon\Gamma\curvearrowright A$ be a cocycle action and let $B\subset(A,\alpha,\mathfrak{w})$. Then there is a map $\mathfrak{u}\colon\Gamma\rightarrow\mathcal{M}(B\otimes\mathbb{K})^{\rm u}$ satisfying the relation $^{\mathfrak{u}}(\mathfrak{w}\otimes 1_{\mathbb{K}})=1$.
@@
We observe that the proof of Lemma 3.1 works in the present situation, after replacing $s$, tsts appearing in the series expansions of $\tilde{c}_{i}$ therein by $s^{\mathfrak{w}}$, $t^{\mathfrak{w}}s^{\mathfrak{w}}$ (not $(ts)^{\mathfrak{w}}$) respectively. Then Lemma 3.3 also works in this setting. By using Lemma 4.2 instead of Lemma 2.2, the proof is complete. ∎
For any simple C-algebra AA, the tensor product $\mathcal{O}_{2}\otimes A$ is purely infinite simple. In particular it contains nonzero projections, and all nonzero projections are mutually Murray–von Neumann equivalent. Moreover, when $A$ is non-unital and separable, for any nonzero $p\in(\mathcal{O}_{2}\otimes A)^{\rm p}$, one can find a sequence $(v_{n})_{n\in\mathbb{N}}$ in $\mathcal{O}_{2}\otimes A$ with
@@ 5. Cocycle actions which are very hard to untwist
satisfying the same axioms as the discrete group case. Here, for a separable C-algebra AA, we equip ${\rm Aut}(A)$, $A$, $\mathcal{M}(A)^{\rm u}$ with the point-norm topology, norm topology, strict topology, respectively. Note that all these topologies are Polish, hence they are standard as Borel spaces. Similarly, we require Borel measurability for the maps witnessing exterior equivalence.
It is fundamental that we only require Borel measurability of the maps. Indeed, for two main motivating examples, namely cocycle actions arising from group extensions and corners of ((cocycle)) actions, the resulting maps are typically discontinuous. However we note that, when $\alpha$ is a homomorphism, it is automatically continuous, because $\operatorname{\mathop{Aut}}(A)$ is a Polish group by [63], Theorem D.11. Similarly, for ordinary actions $\alpha\colon G\curvearrowright A$, any Borel $1$-cocycle $\mathfrak{u}$ of $\alpha$ is automatically strictly continuous. Indeed the map $G\ni s\mapsto\mathfrak{u}_{s}s\in\mathcal{M}(A\mathop{\rtimes_{{\mathrm{r}},\alpha}}G)^{\rm u}$ defines a Borel homomorphism by the cocycle relation, and hence it must be continuous in the strict topology again by [63], Theorem D.11. This proves the strict continuity of $\mathfrak{u}$. As a result, amenability of C-dynamical systems (see [2], [42]) is preserved by exterior equivalence.
@@ 5.1. Corner cocycle actions
For an lcsc group $G$, let $(\alpha,\mathfrak{w})\colon G\curvearrowright A$ be a cocycle action on a separable C-algebra AA. Assume that $p\in A^{\rm p}$ satisfies the following condition: the projections $\alpha_{g}(p)$; $g\in G$ are mutually Murray–von Neumann equivalent. Then one can define the corner cocycle action $(\alpha^{p},\mathfrak{w}^{p})\colon G\curvearrowright pAp$ as follows. We first observe the following lemma.
@@
Let $G$ be a non-exact lcsc group. Then there is a cocycle action $(\alpha,\mathfrak{w})\colon G\curvearrowright\mathcal{O}_{2}$ with the following property. For any cocycle action $(\beta,\mathfrak{x})\colon G\curvearrowright A$ on a unital separable C-algebra, the $(\alpha\otimes\beta)$-22-cocycle $\mathfrak{w}\otimes\mathfrak{x}$ is not a coboundary.
@@
Let $(\beta,\mathfrak{x})\colon G\curvearrowright A$ be any cocycle action on a unital separable C-algebra AA. Suppose that $(\alpha\otimes\beta,\mathfrak{w}\otimes\mathfrak{x})$ is exterior equivalent to a genuine action $\gamma$. Then $(\alpha\otimes\beta,\mathfrak{w}\otimes\mathfrak{x})\otimes 1_{\mathbb{K}}$ is exterior equivalent to $\gamma\otimes 1_{\mathbb{K}}$. By Lemma 5.3, $(\alpha,\mathfrak{w})\otimes 1_{\mathbb{K}}$ is exterior equivalent to $\sigma$. Also $(\beta,\mathfrak{x})\otimes 1_{\mathbb{K}}$ is exterior equivalent to a genuine action by the stabilization trick [55], [44]. Since $\mathbb{K}\cong\mathbb{K}\otimes\mathbb{K}$, these observations show that $\gamma\otimes 1_{\mathbb{K}}$ is exterior equivalent to an amenable action. This implies the amenability of $\gamma$. Since $G$ is non-exact, this contradicts to Corollary 3.6 of [42]. ∎
@@
    When $G$ is exact, such a cocycle action does not exist. Indeed by combining the stabilization trick [55], [44] and the Gabe–Szabo classification theorem [18], one concludes the following statement: Let $\beta\colon G\curvearrowright\mathcal{O}_{2}$ be an amenable, isometrically shift-absorbing action [18]. Then, for any cocycle action $(\alpha,\mathfrak{w})\colon G\curvearrowright A$ on a simple separable nuclear C-algebra (in particular when $A=\mathcal{O}_{2}$), the $(\alpha\otimes\beta)$-22-cocycle $\mathfrak{w}\otimes 1$ is a coboundary. This follows from the proof of Theorem 5.6 of [18]. The proof is similar to the proof of Theorem 6.10 in [18], but for the reader’s convenience, we give a sketch of the proof.
    We first note that when $A$ is non-unital, the claim follows from the stabilization trick (see also Remark 1.6 of [18]). To show the claim in the unital case, we choose an action $\gamma\colon G\curvearrowright A\otimes\mathcal{O}_{2}\otimes\mathbb{K}$ which is exterior equivalent to the stabilization $(\alpha\otimes\beta,\mathfrak{w}\otimes 1)\otimes 1_{\mathbb{K}}$ (which exists by the stabilization trick). Note that for a minimal projection $e\in\mathbb{K}$, the corner cocycle action of $\gamma$ at $p:=1_{A}\otimes 1_{\mathcal{O}_{2}}\otimes e$ is exterior equivalent to $(\alpha\otimes\beta,\mathfrak{w}\otimes 1)$. Let $\eta\colon G\curvearrowright\mathbb{C}$ denote the trivial action. Then by applying the proof of Theorem 5.6 to $0\in{\rm KK}^{G}(\eta,\gamma)$, we obtain a proper cocycle embedding $(\varphi,\mathfrak{v})\colon(\mathbb{C},\eta)\rightarrow(A\otimes\mathcal{O}_{2}\otimes\mathbb{K},\gamma)$ which represents 0 in ${\rm KK}^{G}(\eta,\gamma)$. This implies that the corner cocycle action of $\gamma$ at $q:=\varphi(1)$ is exterior equivalent to a genuine action. Since $q$ is Murray–von Neumann equivalent to $p$, this proves the claim.
@@
    Sutherland [55] proved that any 22-cocycle of a cocycle action on a properly infinite von Neumann algebra is a coboundary. Thus Theorem 5.4 is a very unique phenomenon in C-algebra theory.
@@
6. K0-obstruction: New tensor component $\mathcal{O}_{2}$ is necessary and minimal